OpenStudy (anonymous):

please help!!! Kathleen conducted an experiment in which she drew marbles from a bag one at a time and replaced the marble after each draw. The results of her experiment are shown in the table. Marble Experiment Marble Color Number of Times dawn Red 8 Blue 2 Green 6 Purple 4 Yellow 10According to the results from Kathleen's experiment, what is the probability that she would draw a green marble and then draw a purple marble, if the first marble is replaced before d

OpenStudy (anonymous):

dawning the second marble

OpenStudy (anonymous):

A. 1/90 B. 2/75 C. 3/145 D. 4/145

OpenStudy (anonymous):

i guess from the data we are supposed to conclude that since out of 30 draws she got a green on 6 times then the probability of getting a green one is \(\frac{6}{30}=\frac{1}{5}\) and similarly the probability of getting a purple one is \(\frac{4}{30}=\frac{2}{15}\) multiply these two numbers together to get your answer

OpenStudy (anonymous):

4/30 X 2/15 = 4/225 ?

OpenStudy (anonymous):

6/30 X 1/5 = 1/25 ?

OpenStudy (anonymous):

\[\frac{1}{5}\times \frac{2}{15}=\frac{2}{75}\]